OpenStudy (anonymous):

Let R be the region in the xy-plane between the graphs of y = ex and y = e-x from x = 0 to x = 2. a) Find the volume of the solid generated when R is revolved about the x-axis. (This assignment must be done on graph paper)

OpenStudy (amistre64):

simple enough, you know the formula for the area of a circle?

OpenStudy (anonymous):

yes do i do top - bottom as in \[\int\limits_{0}^{2}\](7.3891-e^x) ^2

OpenStudy (amistre64):

\[\int \pi(e^x)^2dx-\int \pi (e^{-x})^2dx \] \[\int \pi e^{2x}dx-\int \pi e^{-2x}dx \]

OpenStudy (anonymous):

its hallow though. wouldnt you do R^2-r^2

OpenStudy (turingtest):

I have see this problem three times in two days it's like a plague apparently

OpenStudy (turingtest):

seen*

OpenStudy (amistre64):

\[\pi (\frac{e^{2(2)}}{2}-\frac{e^{2(0)}}{2})+ \pi (\frac{e^{-2(2)}}{2}-\frac{e^{-2(0)}}{2}) \] \[\pi (\frac{e^{4}-1}{2})+ \pi (\frac{e^{-4}-1}{2}) \]

OpenStudy (amistre64):

by subtracting the top from the bottom function you create the "hole"

OpenStudy (anonymous):

ok thats what i thought i had to do thanks . so whats the answer

OpenStudy (amistre64):

im sure you are capable enough to finish it up from where I left off ....

OpenStudy (anonymous):

why are you dividing by 2 ?

OpenStudy (amistre64):

integrate e^2x what do you get?

OpenStudy (amistre64):

or better yet derive e^2x and we get 2 e^2x so we need to catch that "2" that pops out by the /2